OpenStudy (anonymous):

(-5, 2) (5, -2) (-5, -2) (5, 2) @kelliegirl33

OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y less than or greater to x - 5 y greater than or equal to -x - 4

OpenStudy (anonymous):

(5,-2) y <= x - 5 -2 <= 5 - 5 -2 <= 0 (true) y >= -x -4 -2 >= -5 - 4 -2 > - 9 (true) (5,-2) is your solution

OpenStudy (anonymous):

ok thank youuu Choose the graph below that represents the following system of inequalities. y greater than or equal to -3x + 1 y less than or greater to 1 over 2x + 3

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

is the second inequality y <= (1/2)x + 3 or y <= 1(2x + 3) ??

OpenStudy (anonymous):

YEEEEEEEES

OpenStudy (anonymous):

the first onw

OpenStudy (anonymous):

it is the 2nd graph

OpenStudy (anonymous):

ok and this is the last one Which of the following points lie in the solution set to the following system of inequalities? y < -3x + 3 y < x + 2 (1, -5) (1, 5) (5, 1) (-1, 5)

OpenStudy (anonymous):

(1,-5) y < -3x + 3 -5 < -3(1) + 3 -5 < -3 + 3 -5 < 0 (true) y < x + 2 -5 < 1 + 2 -5 < 3 (true) solution is (1,-5)